Customers arrive at a checkout counter in a department store according to a Poison distribution of an average of seven per hour.During a given hour,what are the probabilities that

(a).no more than three customers arrive?

(b).at least two customers arrive?

(c).exactly four customers arrive?


1
Expert's answer
2021-05-07T09:23:47-0400

Given, "\lambda=7"

Let X denote the number of customers arrive.


(i) "P(X\le 3)=P(X=1)+P(X=2)+P(X=3)+P(X=4)"


"=\dfrac{e^{-\lambda}\lambda^0}{0!}+\dfrac{e^{-\lambda}\lambda^1}{1!}+\dfrac{e^{-\lambda}\lambda^2}{2!}+\dfrac{e^{-\lambda}\lambda^3}{3!}"


"=\dfrac{e^{-7}7^0}{0!}+\dfrac{e^{-7}7^1}{1!}+\dfrac{e^{-7}7^2}{2!}+\dfrac{e^{-7}7^3}{3!}"


"=0.000911+0.00638+0.0223+0.05212=0.08172"


(ii) "P(X\ge 2)=1-P(X<2)"

"=1-(P(X=0)+P(X=1))"

"=1-(\dfrac{e^{-\lambda}.\lambda^0}{0!}+\dfrac{e^{-\lambda}\lambda^1}{1!})"


"=1-(\dfrac{e^{-7}7^0}{0!}+\dfrac{e^{-7}7^1}{1!})"


"=1-(0.000911+0.0638)=0.9927"


(iii) "P(X=4)=\dfrac{e^{-\lambda}\lambda^4}{4!}"


"=\dfrac{e^{-7}7^4}{24}"


"=0.09122"